# ObjectClear: Complete Object Removal via Object-Effect Attention

### v1

@@ 4 Experiments
Table 1: Quantitative Comparisons with State-of-the-Art Methods. The best and second performances are marked in red and orange, respectively. Although ObjectClear takes only the object mask as input, it outperforms previous methods, even when those methods are provided with masks that cover both the object and its associated effect regions.
@@
Object-Effect Mask SDXL-INP [1] 19.39 22.98 0.2432 0.1024 24.07 26.93 0.1296 0.0579 0.8356
PowerPaint [43] 19.87 21.98 0.2303 0.0776 26.20 27.41 0.1243 0.0409 0.8878
BrushNet [12] 16.82 18.87 0.3434 0.1692 20.96 23.69 0.2052 0.1180 0.6415
DesignEdit [10] 20.69 22.49 0.2946 0.1200 26.59 27.63 0.1777 0.0629 0.9150
CLIPAway [7] 18.87 20.78 0.3328 0.0969 25.38 26.28 0.1039 0.0349 0.8646
FreeCompose [5] 19.67 21.72 0.3316 0.1110 23.39 25.15 0.1305 0.0629 0.8976
Attentive Eraser [31] 20.33 21.98 0.2545 0.1015 27.42 29.27 0.1114 0.0249 0.9095
RORem [15] 21.61 23.11 0.3224 0.0767 27.23 27.95 0.1042 0.0234 0.9157
Object Mask SDXL-INP [1] 20.23 24.83 0.2042 0.0868 22.42 25.77 0.1428 0.0771 0.8044
PowerPaint [43] 21.46 24.62 0.1801 0.0648 22.76 24.67 0.1544 0.0729 0.8538
BrushNet [12] 18.06 23.44 0.2757 0.1821 21.19 24.38 0.1822 0.1123 0.7296
DesignEdit [7] 22.09 24.26 0.2501 0.1021 24.63 25.48 0.1870 0.0788 0.9002
CLIPAway [10] 20.58 23.21 0.2770 0.0785 22.32 24.05 0.1357 0.0765 0.8646
FreeCompose [5] 20.39 22.91 0.3015 0.0897 22.77 24.46 0.1393 0.0690 0.8688
Attentive Eraser [31] 22.17 24.59 0.1883 0.0643 25.70 27.08 0.1201 0.0437 0.9008
RORem [15] 22.49 24.10 0.2943 0.0634 24.51 25.28 0.1288 0.0460 0.9025
ObjectClear (Ours) 26.40 30.75 0.0802 0.0350 32.98 35.14 0.0352 0.0099 0.9172
Implementation. Our proposed method is built on SDXL-Inpainting [1] and fine-tuned with our OBER dataset with an input resolution of $512\times 512$. Training is conducted with a total batch size of $32$ on 8 A100 GPUs for 100k iterations and a learning rate of 1e-5. All experiment results are attained with a guidance scale of $7.5$ and $20$ denoising steps.
Evaluation Data. We evaluate the object removal performance of our method, ObjectClear, against existing approaches on three test datasets: (1) RORD-Val: RORD [29] is a widely used object removal dataset. Each sample contains an image pair (images with and without the target object) and a coarse mask covering both the target object and its associated effect. To avoid duplicated scenarios, we randomly select one image per scene, resulting in a RORD-Val with 343 samples. Since most of the object removal methods, including ObjectClear, require only the object mask as input, we augment RORD-Val with accurate object masks manually to enable a more comprehensive evaluation. (2) OBER-Test: As described in Sec. 3.1, we split our collected OBER dataset into a training set and a test set. The test set contains 160 samples, each comprising image pairs along with precise masks for both the object and its effect. (3) OBER-Wild: To further assess the performance of different methods on out-of-distribution scenarios in the wild, we collect 300 high-quality images featuring objects with associated effects (shadows or reflections) from the Internet. We annotate the object masks using DINO [41] and SAM [13], and manually annotate the effect masks. Note that the OBER-Wild set does not include the removal ground truth.
@@
Figure 4: Object Removal on OBER-Test and RORD-Val. Our ObjectClear effectively removes both the masked objects and their associated effects.
Figure 5: Object Removal on OBER-Wild. Our ObjectClear not only effectively removes single objects with their shadows and reflections (top five samples), but also accurately removes the target object when multiple mutually occluding objects exist (bottom two samples).
@@
Evaluation Metrics. For RORD-Val and OBER-Test, where ground truth is available, we evaluate performance using fidelity metrics: PSNR and PSNR-BG (computed only on background regions), as well as two perceptual metrics: LPIPS [42] and CLIP [25] (feature distance). For OBER-Wild, which lacks ground truth, we employ ReMOVE [3] to measure the consistency between the object removal region and the surrounding background.
Figure 6: Object-Effect Attention Map. ObjectClear achieves a relatively high recall in the object-effect attention map, with recall values increasing as the denoising step progresses. The attention map obtained in the final step effectively covers both the target object and its associated effects, including the object’s shadow (red box) and its reflection in the mirror (green box).
@@ 4.1 Comparisons with State-of-the-Art Methods
We compare ObjectClear with open-source state-of-the-art methods in both image inpainting and object removal. The inpainting methods include SDXL-INP [1], PowerPaint [43], and BrushNet [12]. The object removal methods include CLIP-Away [7], DesignEdit [10], RORem [15], FreeCompose [5], and Attentive Eraser [31]. Following MULAN [32], we use the text prompt “an empty scene” for the text-guided baselines when necessary.
Quantitative Evaluation. Unlike our ObjectClear, which explicitly handles both the target object and its associated effects, most baseline methods operate only within the masked regions, overlooking surrounding areas that are visually affected by the object but not explicitly included in the mask. To ensure a fair and comprehensive comparison, we evaluate all methods under two mask conditions: (1) the input mask covers only the target object, and (2) the input mask covers both the target object and its visual effects. These are referred as Object Mask and Object-Effect Mask settings in Table 1, respectively. As shown in Table 1, ObjectClear achieves state-of-the-art performance across all test sets and metrics. Notably, even when using only the object mask, ObjectClear surpasses methods that rely on both object and effect masks. In particular, it achieves a significant advantage in the PSNR-BG metric, highlighting its superior ability to preserve background consistency with the input.
Qualitative Evaluation. Qualitative results are shown in Fig. 4 and Fig. 5. Generation-based inpainting approaches, such as SDXL-INP [24], BrushNet [12], and PowerPaint [43], often generate new objects within the masked regions while fail to remove the effects of the removed objects. In contrast, previous object removal methods, such as Attentive Eraser [31], DesignEdit [10], and RORem [15] demonstrate strong performance in removing the target object itself, but stil fail to eliminate the associated effects. Thanks to our carefully curated training data and the proposed object-effect attention mechanism, ObjectClear effectively removes not only the target object but also its associated shadows and reflections. Moreover, as shown in the last two rows of Fig. 5, ObjectClear effectively handles complex cases involving multiple mutually occluding objects.
Table 2: Quantitative Results of Ablation Studies. Based on our camera-captured data (CC Data), our object removal performance improves progressively by adding the object-effect mask loss ($\mathcal{L}_{mask}$) and simulated training data (Sim. Data). Further improvement is achieved through the proposed Attention-Guided Fusion (AG Fusion).
@@
(a) 27.29 27.96 0.0910 0.0247
(b) 27.56 28.37 0.0845 0.0217
(c) 28.00 28.81 0.0818 0.0190
(d) 32.98 35.14 0.0352 0.0099
@@ 4.2 Ablation Studies
The Effectiveness of the Object-Effect Mask Loss $\mathcal{L}_{mask}$. The object-effect mask loss is designed to explicitly guide ObjectClear to attend to both the target object and its associated effect regions, leading to more thorough and natural removal results. As shown in the chart on the left of Fig. 6, the predicted object-effect attention map achieves a high recall of 0.97. Besides, the attention map in the last column of Fig. 6 on the right, which is from the final denoising step (Step 20), effectively reveals not only the object’s shadow (marked with a red box), but also the object’s reflection in the mirror (marked with a green box), altogether leading to complete and visually coherent object removal results. From Table 2 (a) to (b), improvements can be seen across all metrics with $\mathcal{L}_{mask}$ added, indicating its effectiveness in improving object removal accuracy.
The Effectiveness of Simulated Data. The simulated data effectively enhances the background diversity and scenario complexity of the training set, leading to notable improvements in object removal performance and better background preserving, as shown in Table 2(c). Our carefully designed simulation pipeline, especially the multi-object data synthesis, enables ObjectClear to handle challenging cases involving multiple mutually occluding objects, as demonstrated in the last two rows of Fig. 5. More visualized results can be found in the appendix.
Effectiveness of the Attention-Guided Fusion. The attention-guided fusion operation is designed to preserve background consistency during object removal, a common challenge in generation-based methods such as ObjectDrop [34] and RORem [15], as illustrated in Fig. 7. As the denoising process progresses, the proposed object-effect attention yields increasingly accurate attention maps (Fig.6), which we leverage to guide the fusion of background images. This strategy significantly improves background preservation, as evidenced by notable gains in PSNR-BG. Consequently, we observe a substantial boost in overall performance, as reflected in the quantitative results from (c) to (d) in Table 2, where PSNR improves markedly from $28.00$ to $32.98$.
Figure 7: Effectiveness of Attention-Guided Fusion. The difference maps visualize the pixel-wise differences between the input and output images. RORem [15] fails to remove the reflection of the balloon and introduces changes to the background, resulting in relatively high difference values in the background area and low values in the reflection region. In contrast, both ObjectDrop [34] and our ObjectClear successfully remove the reflection. Notably, ObjectClear better preserves the background, thanks to our predicted object-effect map and proposed attention-guided fusion strategy.
4.3 More Applications
ObjectClear can be flexibly extended to various applications, such as object insertion and movement. We also utilize our OBER dataset to train an insertion network that is able to generate natural object effects. Please refer to the appendix for more details. As shown in Fig. 8, even when only the target objects are specified for insertion and movement, ObjectClear is capable of generating plausible and natural shadows and reflections accordingly.
@@ 5 Conclusion
We introduce ObjectClear, a practical framework for object removal that ensures high-quality object-effect removal and background-preserving results across diverse real-world scenarios. Our method employs an Object-Effect Attention (OEA) mechanism to adaptively focus on removal regions, including both the object and its effects, coupled with Attention-Guided Fusion (AGF) to preserve background details accurately. We also propose OBER, a large-scale and diverse hybrid dataset combining camera-captured and simulated data. With the carefully designed mechanism and curated dataset, ObjectClear achieves robust and superior performance, effectively addressing existing challenges in object removal tasks. In addition, ObjectClear can be easily adapted to various applications, such as object insertion and object movement, and consistently achieves superior performance. These advancements position ObjectClear as a reliable solution for practical image editing, laying a solid foundation for future research in controllable image processing through local editable regions.
@@
In this supplementary material, we provide additional discussions and results to complete the main paper. In Sec. A, we present further details on data augmentation during training and other extended applications. In Sec. B, we provide more details of the proposed OBER dataset, including statistics and some examples for demonstration. In Sec. C, we report additional results, such as more results on ablation studies, a user study, experiments with user strokes, more results of object insertion and movement, and additional comparisons on in-the-wild data.
A More Details of ObjectClear
A.1 Training Augmentations
@@
During training, we apply on-the-fly random cropping to enable the network to learn object removal across varying object sizes. We also apply color augmentation and random flipping to enhance model robustness. To improve the model’s robustness to the estimated or user-provided coarse mask, several previous methods [34, 11, 33] have introduced mask augmentation techniques. In line with these approaches, we also apply dilation and erosion to the input mask during training. Different from previous practice, our method employs an object-aware dilation and erosion strategy, where the dilation and erosion kernel size is adaptively determined based on the size of the object. As demonstrated in the qualitative results in Sec. C.3, our method effectively handles coarse user-drawn masks by implicitly completing and refining them, showcasing strong robustness to imprecise inputs.
A.2 More Details for Other Applications
In this section, we further discuss implementation details of other applications, including object insertion and object movement.
Object Insertion. The insertion network also leverages the OBER dataset for training and adopts the same architecture as ObjectClear (Fig. 3 of the main paper), thus receiving the input tuple of $<z_{t},I_{in},M_{o},c>$ and supervised by $I_{GT}$ (for output image) and $M_{fg}$ (for object-effect attention). However, $I_{in}$, $I_{GT}$, and $c$ are constructed in an “reverse” way compared with the removal network.
@@ A.2 More Details for Other Applications
While the ground truth $I_{GT}$ is the original image with the object accompanied by natural effects, $I_{in}$ is the one with the object simply copied and pasted on the background. Specifically, to generate $I_{in}$, we first extract the object from $I_{GT}$ with the object mask $M_{o}$ only, and then we paste it onto the corresponding background image. In addition to the image pair, the input text $c$ is also reversed to “insert the instance of”. Notably, the object-effect attention map is still supervised by $M_{fg}$. In this context, $M_{fg}$ refers to the mask that covers both the object and its generated effects (e.g., shadows or reflections).
@@
During inference, we directly paste an object (without effects) onto the background scene, along with its object mask $M_{o}$, which forms the input image $I_{in}$. The network then generates the output image where both the object and its generated effects are harmoniously inserted. Since the insertion network also integrates the object-effect attention, we can also apply Attention-Guided Fusion to preserve the background fidelity while generating realistic object effects.
@@
Object Movement. To enable object movement, we combine the ObjectClear and insertion network introduced above. Specifically, we first apply ObjectClear to remove the target object along with its associated effects (e.g., shadows or reflections), resulting in a clean object-free background. The object is then extracted using its given object mask. Users are allowed to specify a new location and optionally adjust the object scale before the object is re-inserted into the clean background. With our insertion network, the object is harmonized with the new context by generating realistic effects. This two-stage approach supports controllable object movement while ensuring visual realism and consistency.
B More Details of OBER Dataset

### v2

@@ 4 Experiments
Table 1: Quantitative Comparisons with State-of-the-Art Methods. The best and second performances are marked in red and orange, respectively. Although ObjectClear takes only the object mask as input, it outperforms prior methods on most metrics, even when some methods are provided with masks that cover both the object and its associated effect regions.
@@
Object-Effect Mask SDXL-INP [26] 19.39 22.98 0.2432 0.1024 24.07 26.93 0.1296 0.0579 0.6983
PowerPaint [48] 19.87 21.98 0.2303 0.0776 26.20 27.41 0.1243 0.0409 0.8044
BrushNet [12] 16.82 18.87 0.3434 0.1692 20.96 23.69 0.2052 0.1180 0.5358
DesignEdit [10] 20.69 22.49 0.2946 0.1200 26.59 27.63 0.1777 0.0629 0.8215
CLIPAway [6] 18.87 20.78 0.3328 0.0969 25.38 26.28 0.1039 0.0349 0.7705
FreeCompose [4] 19.67 21.72 0.3316 0.1110 23.39 25.15 0.1305 0.0629 0.7555
Attentive Eraser [34] 20.33 21.98 0.2545 0.1015 27.42 29.27 0.1114 0.0249 0.7940
RORem [18] 21.61 23.11 0.3224 0.0767 27.23 27.95 0.1042 0.0234 0.8164
Object Mask SDXL-INP [26] 20.23 24.83 0.2042 0.0868 22.42 25.77 0.1428 0.0771 0.6971
PowerPaint [48] 21.46 24.62 0.1801 0.0648 22.76 24.67 0.1544 0.0729 0.7699
BrushNet [12] 18.06 23.44 0.2757 0.1821 21.19 24.38 0.1822 0.1123 0.6341
DesignEdit [10] 22.09 24.26 0.2501 0.1021 24.63 25.48 0.1870 0.0788 0.8163
CLIPAway [6] 20.58 23.21 0.2770 0.0785 22.32 24.05 0.1357 0.0765 0.7705
FreeCompose [4] 20.39 22.91 0.3015 0.0897 22.77 24.46 0.1393 0.0690 0.7451
Attentive Eraser [34] 22.17 24.59 0.1883 0.0643 25.70 27.08 0.1201 0.0437 0.8080
RORem [18] 22.49 24.10 0.2943 0.0634 24.51 25.28 0.1288 0.0460 0.8121
OmniEraser [37] 21.79 22.98 0.2195 0.0542 24.44 24.87 0.1783 0.0142 0.7655
GeoRemover [47] 24.38 26.47 0.1924 0.0575 24.92 25.33 0.1976 0.0670 0.7938
OmniPaint [42] 22.75 24.66 0.1178 0.0354 29.06 30.04 0.0521 0.0152 0.8656
ObjectClear (Ours) 26.24 29.78 0.1157 0.0299 33.04 35.62 0.0342 0.0103 0.8470
@@
Figure 4: Object Removal on OBER-Test and RORD-Val. Our ObjectClear effectively removes both the masked objects and their associated effects, including shadows and mirror reflections.
Implementation. Our network is built on SDXL-Inpainting [26] and fine-tuned with our OBER dataset with an input resolution of $512\times 512$. Training is conducted with a total batch size of $32$ on 8 A100 GPUs for 100k iterations and a learning rate of 1e-5. All experiment results are attained with a guidance scale of $1.0$ and $20$ denoising steps.
Figure 5: Object Removal on OBER-Wild. Prior methods may not completely remove object effects and can also introduce residual artifacts or unwanted content. In contrast, ObjectClear achieves precise object–effect removal while preserving high background fidelity.
Evaluation Data. We evaluate ObjectClear and prior methods on three test sets: (1) RORD-Val. RORD [31] is a widely used object removal dataset. Each sample contains an image pair (w/ and w/o object) and a coarse mask covering both object and its effect. To avoid duplicated scenarios, we randomly select one image per scene, resulting in a RORD-Val with 343 samples. Since some object removal methods, including OmniEraser [37], OmniPaint [42] and ours, require only the object mask as input, we manually refine the object masks to ensure fair evaluation. (2) OBER-Test. As described in Sec. 3.1, we split our OBER dataset into training and test sets. The test set contains 163 samples, each providing image pairs and precise masks for both object and effect. (3) OBER-Wild. To assess generalization in real-world scenarios, we collect 302 in-the-wild images featuring objects with shadows or reflections. Object masks are annotated using DINO [44] and SAM [13], while effect masks are manually labeled. Ground-truth removal images are not available for this set.
@@
Evaluation Metrics. For RORD-Val and OBER-Test, where ground truth is available, we evaluate performance using fidelity metrics: PSNR and PSNR-BG (computed only on background regions), as well as two perceptual metrics: LPIPS [45] and CLIP [27] (feature distance). For OBER-Wild, which lacks ground truth, we evaluate results using ReMOVE, a modified version of ReMOVE [2] designed to better assess visual harmony. Specifically, ReMOVE measures the consistency between the output’s removal region and the input background, instead of comparing against the output background as in the original ReMOVE.
@@ 4.1 Comparisons with State-of-the-Art Methods
We compare ObjectClear with state-of-the-art methods, including: 1) image inpainting methods of SDXL-INP [26], PowerPaint [48], and BrushNet [12]; 2) object removal methods of CLIP-Away [6], DesignEdit [10], RORem [18], FreeCompose [4], Attentive Eraser [34], OmniEraser [37], GeoRemover [47], and OmniPaint [42].
Quantitative Evaluation. Unlike our ObjectClear, which explicitly handles both the target object and its associated effects, most baseline methods operate only within the masked regions, overlooking surrounding areas that are visually affected by the object but not explicitly included in the mask. To ensure a fair and comprehensive comparison, we test all methods under two mask conditions: (1) Object Mask, and (2) Object-Effect Mask, as reported in Table 1. ObjectClear achieves state-of-the-art performance on nearly all test sets and metrics. Notably, even when using only the object mask, ObjectClear surpasses methods that rely on both object and effect masks. In particular, it achieves a significant advantage in the PSNR-BG metric, highlighting its superior ability to preserve background consistency with the input.
Table 2: Quantitative Results of Ablation Studies. Based on our camera-captured data ($C$ Data), object removal performance improves progressively by introducing the Adaptive Target-Aware Attention (ATA), simulated training data (Sim. Data), Attention-Guided Fusion (AGF), and Spatially-Varying Denoising Strength (SVDS) strategy.
@@
(a) 0.99 27.29 27.96 0.0910 0.0247
(b) 0.99 27.56 28.37 0.0845 0.0217
(c) 0.99 28.04 28.80 0.0805 0.0196
(d) 0.99 32.77 35.50 0.0348 0.0106
(e) 1.00 31.49 33.46 0.0375 0.0120
(f) SVDS 33.04 35.62 0.0342 0.0103
Figure 6: Object-Effect Map from Adaptive Target-Aware Attention. ObjectClear achieves a relatively high recall in the object-effect map, with recall values increasing as the denoising step progresses. The attention map obtained in the final step effectively covers both the target object and its associated effects, including the object’s shadow (red box) and its reflection in the mirror (green box).
Figure 7: Effectiveness of Attention-Guided Fusion. The difference maps visualize the pixel-wise differences between input and output. RORem [18] fails to remove the reflection and alters the background, yielding low differences in the reflection area and high values in the background. In contrast, both ObjectDrop [38] and our ObjectClear successfully remove the reflection, with ObjectClear better preserving the background as indicated by lower differences.
Qualitative Evaluation. Qualitative results are shown in Fig. 4 and Fig. 5. Generation-based inpainting approaches, such as SDXL-INP [26], and PowerPaint [48], often generate new objects within the masked regions while fail to remove the effects of the removed objects. In contrast, previous object removal methods, such as Attentive Eraser [34], DesignEdit [10], and RORem [18] demonstrate strong performance in removing the target object itself, but still fail to eliminate the associated effects. OmniEraser [37] demonstrates the ability to remove both objects and their effects, though it may still leave residual objects or effects, occasionally erase non-target regions, or introduce background changes. Similarly, OmniPaint [42] maintains good background fidelity but may occasionally fail to fully remove the target object and its associated effects. Benefiting from our training data and network designs, ObjectClear excels at accurately removing both the target object and its effects. Moreover, it robustly handles complex scenarios involving multiple occluding objects (see the last row of Fig. 5).
@@ 4.2 Ablation Studies
The Effectiveness of Adaptive Target-Aware Attention (ATA). The ATA is designed to explicitly guide ObjectClear to attend to both the target object and its effects, leading to more complete and precise removal. From Table 2 (a) to (b), improvements can be seen across all metrics with ATA introduced, indicating its effectiveness in improving object removal accuracy (see visual comparison in supp). Besides, as shown in the plot of Fig. 6 (left), the object-effect map from ATA module achieves a high recall of 0.97. In Fig. 6 (right), the attention map from the final denoising step (Step 20) effectively reveals object’s shadow (red box) and mirror reflection (green box), altogether leading to complete and visually coherent object removal results.
The Effectiveness of Simulated Data. The simulated data effectively enhances the scale and diversity of the training data, leading to notable improvements in object removal performance and better background preserving, as shown in Table 2(c). Our carefully designed simulation pipeline, especially the multi-object data synthesis, further enables ObjectClear to tackle challenging scenarios involving multiple objects. We provide more visual results in supplementary.
Effectiveness of the Attention-Guided Fusion (AGF). The AGF is designed to preserve background consistency after object removal, a common challenge in generation-based methods like ObjectDrop [38] and RORem [18], as illustrated in Fig. 7. As the denoising process, our Adaptive Target-Aware Attention yields increasingly accurate attention maps (Fig.6), which we leverage to guide the fusion of input background. This strategy significantly improves background preservation, as evidenced by large gains in PSNR-BG. Consequently, we observe a marked boost in overall performance, with PSNR increasing from $28.04$ to $32.77$, as shown in Table 2(c–d).
Effectiveness of the Spatial-Varying Denoising Strength. The proposed Spatially-Varying Denoising Strength (SVDS) addresses two major issues: incomplete object removal or hallucinated objects when using $\mathrm{DS}=0.99$, and global color shifts when using $\mathrm{DS}=1.0$ (see visual ablation in supp.). As shown in Table 2(d–f), SVDS (f) outperforms both $\mathrm{DS}=0.99$ (d) and $\mathrm{DS}=1.0$ (e). It effectively facilitates complete object removal while maintaining consistent background colors and details.
@@ 5 Conclusion
We introduce ObjectClear, a practical framework for object removal that achieves high-quality object-effect removal while maintaining background consistency across diverse real-world scenarios. The framework employs an Adaptive Target-Aware Attention to adaptively focus on removal regions, along with Attention-Guided Fusion and Spatial-Varying Denoising Strength strategy to preserve background details. In addition, we present OBER, a large-scale and diverse hybrid dataset that integrates camera-captured and simulated data. Thanks to our dataset and network designs, ObjectClear shows robust and superior performance, effectively overcoming key challenges in object removal and laying a solid foundation for future research in this field.
Acknowledgement. This research is supported by cash and in-kind funding from NTU S-Lab and industry partner(s).
@@
In this appendix, we provide additional discussions and results to complete the main paper. In Sec. 6, we present further implementation details on data augmentation during training and other extended applications. In Sec. 7, we provide a more detailed introduction to the proposed OBER dataset, including statistics and some examples for demonstration. In Sec. 8, we report additional results, such as further ablations, a user study, experiments with user strokes, results for object insertion and movement, and additional comparisons on in-the-wild images. In addition, we provide an online [interactive demo] on Hugging Face, enabling users to remove arbitrary objects through simple clicks.
6 Implementation Details
6.1 Training Augmentations
@@
During training, we apply on-the-fly random cropping to enable the network to learn object removal across varying object sizes. We also apply color augmentation and random flipping to enhance model robustness. To improve the model’s robustness to the estimated or user-provided coarse mask, several previous methods [38, 11, 37] have introduced mask augmentation techniques. In line with these approaches, we also apply dilation and erosion to the input mask during training. Different from previous practice, our method employs an object-aware dilation and erosion strategy, where the dilation and erosion kernel size is adaptively determined based on the size of the object. As demonstrated in the qualitative results in Sec. 8.6, our method effectively handles coarse user-drawn masks by implicitly completing and refining them, showcasing strong robustness to imprecise inputs.
6.2 More Details for Extended Applications
ObjectClear can be flexibly extended to various applications. Visual results of object insertion and movement are shown in Fig. 16, where our models generate realistic visual effects. In this section, we provide the implementation details of object insertion and object movement.
@@ 6.2 More Details for Extended Applications
Object Insertion. The insertion network also leverages the OBER dataset for training and adopts the same architecture as ObjectClear, as illustrated in Fig.3 of the main paper, which receives the input tuple of $z_{t},I_{in},M_{o},c>$ and supervised by $I_{GT}$ (for output image) and $M_{fg}$ (for Adaptive Target-Aware Attention). However, $I_{in}$, $I_{GT}$, and $c$ are constructed in a reverse manner compared with the removal network. Specifically, the ground-truth image $I_{GT}$ corresponds to the original image containing the object along with its associated effects, while $I_{in}$ is obtained by simply copying and pasting the object onto the background without any effects. To generate $I_{in}$, we first extract the object from $I_{GT}$ using the object mask $M_{o}$ only, and then paste it into the corresponding background image. In addition to the image pair, the input text $c$ is also reversed to “insert the instance of”. Notably, the Adaptive Target-Aware Attention map is still supervised by $M_{fg}$. Here, $M_{fg}$ refers to the mask that covers both the object and its generated effects (e.g., shadows or reflections).
@@
During inference, we obtain the input image $I_{in}$ by pasting an object foreground (w/o effects) onto the background image, and we feed it together with its corresponding object mask $M_{o}$ as the input pair. The network then generates the output image where both the object and its generated effects are harmoniously inserted. Since the insertion network also integrates the Adaptive Target-Aware Attention, we further apply the Attention-Guided Fusion to preserve background fidelity while synthesizing realistic object effects.
@@
Object Movement. To enable object movement, we combine our object removal (ObjectClear) and insertion models into a two-stage framework. Specifically, we first apply ObjectClear to remove the target object and its associated effects, producing a clean, object-free background. The object is then extracted using its provided object mask, and users can specify a new location and adjust its scale before reinsertion. With our insertion network, the object is harmonized with the new context by generating realistic effects. Without retraining any additional networks, this two-stage pipeline supports controllable object movement while maintaining visual realism and consistency.
7 More Details of OBER Dataset
Figure 8: Samples from our OBER Dataset. We provide some samples from our OBER dataset including both camera-captured data and simulated data. While all data from both categories are annotated with fine-grained object masks and object-effect masks, we also extract RGBA foreground objects with associated effects from camera-captured data (as discussed in Sec. 3.1 in the main paper), which could be used to construct realistic simulated samples. In the simulated samples, we not only include the shadow effects but also the reflection effects, which are rarely included in other datasets, thanks to our reflection pair simulation strategy detailed in Sec. 7.